OpenStudy (anonymous):

Hitomi has a bag with four chips in it numbered 1 to 4. She randomly draws a chip and replaces it three times. What is the probability that Hitomi will draw a 1 all three times?

OpenStudy (anonymous):

IT HA TO BE A FRaCTION @Bookworm14

OpenStudy (bookworm14):

sorry girly, im not sure about this one :/, i know it, i just cant think of it @EclipsedStar @Teddyiswatshecallsme @bohotness @TheSmartOne @sleepyjess

OpenStudy (anonymous):

okay, thx anyhow

OpenStudy (anonymous):

anyone else? plz!

TheSmartOne (thesmartone):

If she draws a 1 the first time, that is a chance of 1/4 The second time again drawing a 1 is another 1/4 And the same with the third time. So what is \(\sf\Large\frac{1}{4}\times\frac{1}{4}\times\frac{1}{4}=?\)

OpenStudy (bohotness):

YES

OpenStudy (bookworm14):

@TheSmartone @bohotness thank you for helping her :)

TheSmartOne (thesmartone):

Anytime :D